x+5y+3z=4
4y+2z=6
z=-3
solve

Answer by checkley71(8403)

Z=-3 THEN 4Y+2(-3)=6 OR 4Y-6=6 OR 4Y=12 OR Y=12/4 OR Y=3 THEN X+5*3+3*(-3)=4 OR
X+15-9=4 OR X+6=4 OR X=4-6 OR X=-2